OpenStudy (anonymous):

A bag contains 5 red marbles, 6 white marbles, and 5 blue marbles. Find p(red and blue)

OpenStudy (anonymous):

the probability is 10/16 reduced to 5/8

OpenStudy (kirbykirby):

^wouldn't that be P(red OR blue)

OpenStudy (anonymous):

5/8 is what i got for P(red OR blue)

OpenStudy (anonymous):

no because the question says red "AND" blue @kirbykirby

OpenStudy (kirbykirby):

1st draw: getting a red is 5/16 2nd draw: getting a blue is 5/15, since there are 15 marbles left. Then you use the product rule

OpenStudy (anonymous):

no.....just no

OpenStudy (anonymous):

I'm so confused

OpenStudy (kirbykirby):

Then how would you find P(red and blue and white and red) ?

OpenStudy (kirbykirby):

If you do by adding the individual probabilities, then you will go over 1 and probabilities can't exceed 1

OpenStudy (anonymous):

"I'M DONE" Because me and @tmason8 are correct so it's over

OpenStudy (turingtest):

your problem doesn't specify whether this is a single draw of two marbles, or two draws, or whether marbles are put back.... please clarify, or the question cannot be answered

OpenStudy (anonymous):

huh?

OpenStudy (turingtest):

if the marbles are drawn successively and not put back, then @kirbykirby is correct I have no idea where you got your answer from @gswag98

OpenStudy (anonymous):

The question doesn't specify that, which is why I am so confused. I am used to them saying if it is single draw etc.

OpenStudy (turingtest):

a single draw is the same as tow successive draws without any marbles being put back. As @kirbykirby stated, P(red)=5/16, then P(blue)=4/15 since once less marble is now in the bag. You can then treat the two events as independent.

OpenStudy (turingtest):

typo P(blue)=5/15

OpenStudy (kirbykirby):

Now that I think of it, if it's not successive draws, then we'd have to divide by 2 right? or think of it as \[ \large \frac{{5 \choose 1}{5 \choose 1}}{{16 \choose 2}}\]

OpenStudy (kirbykirby):

*multiply by 2

OpenStudy (turingtest):

no, i don't think it would change your answer if the draws are not successive, because\[P(R\cup B)=P(R)P(B|R)=P(B)P(R|B)\]

OpenStudy (turingtest):

the conditional probabilities for red/blue respectively are the same regardless of order, and drawing 'at the same time' can always be represented as successive draws without replacement

OpenStudy (anonymous):

What is the mean, variance, and standard deviation of the values? Round to the nearest tenth. 1, 9, 4, 12, 13, 13

OpenStudy (turingtest):

please post each question separately

OpenStudy (anonymous):

how?

OpenStudy (turingtest):

close this post (little blue button near start of thread) then do a new one just like before

OpenStudy (kirbykirby):

Hm but if we draw blue first then red, the order doesn't matter, so are we double counting? @TuringTest

OpenStudy (turingtest):

hmm i think you are right

OpenStudy (turingtest):

by the total probability theorem, i think\[P(R\cup B)=P(R)(P(B|R)+P(B)P(R|B)\]right? so that makes you right, we need to multiply your original answer by 2 since the terms above are equal man, i just took probability and i'm already forgetting it :P

OpenStudy (turingtest):

yes of course you are right, silly me XD

OpenStudy (anonymous):

don't make sense

OpenStudy (turingtest):

we went above your level, I think. sorry, but you probably have a set way to do this from your book, but kirby and I are going off general probability laws

OpenStudy (anonymous):

ok

OpenStudy (kirbykirby):

Ah yes total probability rule hehe.

OpenStudy (anonymous):

@LearningIsAwesome

OpenStudy (anonymous):

Can you make a new question and tag me in it? I think I know this

OpenStudy (anonymous):

Can you make a new question and tag me in it? I think I know this @tmason8

OpenStudy (kirbykirby):

\[\large \frac{{5 \choose 1}{5 \choose 1}}{{16 \choose 2}} \] Or think of P(red and blue) as [P(red then blue) or P(blue then red] = P(red then blue) + P(blue then red) = (5/16)(5/15) + (5/16)(5/15)

OpenStudy (anonymous):

hmmm... why is this still being replied to I mean tired of getting a notification of this problem you know I think me and @tmason8 are right that's all too it

OpenStudy (kirbykirby):

I know that you can add the fractions 5/16 + 5/16 together

OpenStudy (kirbykirby):

can't*

OpenStudy (kirbykirby):

that's for P(red "or" blue)